### v1

# Pessimistic Iterative Planning for Robust POMDPs

Abstract
Robust partially observable Markov decision processes (robust POMDPs) extend classical POMDPs to handle additional uncertainty on the transition and observation probabilities via so-called uncertainty sets. Policies for robust POMDPs must not only be memory-based to account for partial observability but also robust against model uncertainty to account for the worst-case instances from the uncertainty sets. We propose the pessimistic iterative planning (PIP) framework, which finds robust memory-based policies for robust POMDPs. PIP alternates between two main steps: (1) selecting an adversarial (non-robust) POMDP via worst-case probability instances from the uncertainty sets; and (2) computing a finite-state controller (FSC) for this adversarial POMDP. We evaluate the performance of this FSC on the original robust POMDP and use this evaluation in step (1) to select the next adversarial POMDP. Within PIP, we propose the rFSCNet algorithm. In each iteration, rFSCNet finds an FSC through a recurrent neural network trained using supervision policies optimized for the adversarial POMDP. The empirical evaluation in four benchmark environments showcases improved robustness against a baseline method in an ablation study and competitive performance compared to a state-of-the-art robust POMDP solver.
1 Introduction
Partially observable Markov decision processes (POMDPs; Kaelbling et al.,, 1998) are the standard model for decision-making under uncertainty. Policies select actions based on limited state information towards some objective, e.g., minimizing the expected cost. The standard assumption is that probabilities are precisely known for the transition and observation functions of a POMDP. This assumption is unrealistic, for example, when probabilities are derived from historical data or sensors with limited precision (Thrun et al.,, 2005), or when the uncertainty is expressed by domain experts.
Robust POMDPs (RPOMDPs; Osogami,, 2015) overcome this assumption by introducing uncertainty sets, i.e., sets of probabilities, on the transition and observation functions. Robust policies for RPOMDPs account for this uncertainty by optimizing against the worst-case instances within the uncertainty sets. As a consequence, a robust policy has a lower bound on its actual performance.
Finding optimal policies for (non-robust) POMDPs is an extensively studied problem (Poupart and Boutilier,, 2003; Hansen,, 1997; Smith and Simmons,, 2004). The policies inherently depend on the sequences of past actions and observations (the history), i.e., they require memory. Recurrent neural networks (RNNs) are a common formalism to represent such memory-based policies (Ni et al.,, 2022) thanks to their ability to learn sufficient statistics of the history (Lambrechts et al.,, 2022). As such, RNNs have successfully been used in POMDPs within reinforcement learning (Bakker,, 2001; Hausknecht and Stone,, 2015; Heess et al.,, 2015) or planning settings (Carr et al.,, 2019, 2020, 2021).
Alternatively, finite-state controllers (FSCs, Meuleau et al.,, 1999) offer a more structured way to represent policies and, given the model, allow for a precise and effective numerical evaluation of their performance. Finding FSCs often relies on selecting a predetermined memory size and structure (Junges et al.,, 2018). An exhaustive search for the optimal size and specific structures may, in general, be computationally intractable. Therefore, one may decide to cluster any history-based policy into a finite-memory policy, for instance, by expanding all possible histories and minimizing the resulting so-called policy tree (Grzes et al.,, 2015). However, such an approach suffers from the typical exponential blow-up in the size of the history that is inherent to POMDPs. Instead, Carr et al., (2021) learn the memory structures for FSCs by training an RNN on data collected from the POMDP.
Challenges in RPOMDPs. Finding robust policies for RPOMDPs has its own set of challenges. As robust policies need to be optimized against the worst-case instance of the uncertainty sets, determining this worst-case is paramount. Furthermore, a policy’s performance in an RPOMDP cannot be evaluated through simulations, as it is unclear which instance from the uncertainty sets to sample from. This last challenge is detrimental to the use of RNNs in planning settings where guarantees on a policy’s worst-case performance are required. FSCs, on the other hand, can be evaluated exactly in RPOMDPs through robust dynamic programming (Iyengar,, 2005), which finds and uses the exact worst-case instance to evaluate the FSC’s performance. Finding robust policies represented as FSCs, however, suffers from the same problems as in POMDPs, e.g., selecting a predetermined memory size and structure, as done in sequential convex programming (SCP, Cubuktepe et al.,, 2021). As our empirical evaluation shows, besides being computationally intractable to establish the optimal FSC size and structure, wrongly or over-specifying the required memory may lead to worse results in RPOMDPs.
(1) Supervised learning of robust FSCs (2) Adversarial POMDP selection Robust dynamic programming RNN $\rightarrow$ FSC POMDP RNN FSC Robust value
Figure 1: Overview of the PIP framework. Steps on the left-hand side are specific to rFSCNet.
Contributions. Observing how RPOMDPs, RNNs, and FSCs complement each other, we combine the representational power of RNNs and the exact robust evaluation of FSCs, allowing for the best of both worlds by leveraging prior work to derive an FSC representation of an RNN policy (Koul et al.,, 2019; Carr et al.,, 2021).
As outlined in Figure 1, our main contribution is an iterative pessimistic planning framework.
  • ($\star$)
    Pessimistic iterative planning for RPOMDPs. We propose PIP, a framework to find a robust policy by iterating over and (approximately) solving POMDPs within the uncertainty set of the RPOMDP that are adversarial instances to the current policy, i.e., a worst-case instance given the current policy. We implement PIP in our RNN-based algorithm, which we name rFSCNet, consisting of the following two main parts:
  1. 1.
    Supervised learning of robust FSCs (Section 4). We train an RNN based on data collected from supervision policies that we compute on the adversarial POMDPs. We explicitly construct the RNN to enable the extraction of FSCs. At each iteration, the FSC serves in the robust evaluation and adversarial POMDP selection. The adversarial instances are employed to further train the RNN by refining the collected histories and associated supervision policy, guiding the RNN towards a better-performing, robust policy.
  2. 2.
    Adversarial POMDP selection (Section 5). First, we compute the worst-case performance of the FSC on the RPOMDP via robust dynamic programming, thereby producing an exact guarantee. Then, we introduce a linear program that efficiently finds a POMDP instance within the uncertainty sets that is adversarial, i.e., a worst-case instance, to the current FSC.
In our experimental evaluation on four benchmarks, we (1) showcase that the FSCs found by our method are competitive with and outperform the state-of-the-art FSC-based solver for RPOMDPs by Cubuktepe et al., (2021), and, (2) conduct an ablation study to show the impact of our contributions on robustness, namely iteratively finding and re-training on adversarial POMDPs, as opposed to a baseline that trains on a fixed POMDP within the uncertainty set of the RPOMDP.
2 Preliminaries
The set of all distributions over $X$ is denoted by $\Delta(X)$. A probability distribution $\mu\in\Delta(X)$ is called Dirac if $\mu(x)=1$ for precisely one $x\in X$ and zero otherwise. The number of elements in a set $X$ is denoted by $|X|$. Iverson brackets return $[P]=1$ if $P$ is true and 00 otherwise. Finally, the set of probability intervals with lower bounds strictly greater than zero is $\mathbb{I}=\{[i,j]\mid 0<i\leq j\leq 1\}$.
Definition 1 (POMDP).
A partially observable Markov decision process is a tuple $M=\braket{S,A,T,Z,O,C}$, where $S,A,Z$ are finite sets of states, actions, and observations, $T\colon S\times A\to\Delta(S)$ is the probabilistic transition function, $O\colon S\to Z$ is the deterministic state-based observation function, and $C\colon S\times A\to\mathbb{R}_{\geq 0}$ is the cost (or reward) function.
For simplicity and without loss of generality, we consider POMDPs with deterministic observations, where the observation function maps only to Dirac distributions. This assumption is non-restrictive as every POMDP can be transformed into one with deterministic observations (Chatterjee et al.,, 2015, Remark 1). A trajectory in a POMDP is a finite sequence of states and actions: $\omega_{t}=s_{1}a_{1}s_{2}\dots s_{t}\in(S\times A)^{t-1}\times S$, such that $T(s_{i+1}\mid s_{i},a_{i})>0$ for $0\leq i<t$. Due to partial observability, states are not observable to the agent. A history is the observable part of a trajectory: $h_{t}=O(s_{1})a_{1}O(s_{2})\dots O(s_{t})=z_{1}a_{1}z_{2}\dots z_{t}\in(Z\times A)^{t-1}\times Z$. The set of all (finite) trajectories and associated histories are $\Omega$ and $\mathcal{H}$, respectively. Histories can be compressed into sufficient statistics known as beliefs, that is, probability distributions over states (Kaelbling et al.,, 1998). The set of all belief states in a POMDP is $\mathcal{B}\subseteq\Delta(S)$. The initial state distribution (belief) is $b_{0}\in\mathcal{B}$. A belief $b\in\mathcal{B}$ can be computed from a history $h\in\mathcal{H}$ using Bayes’ rule (Spaan,, 2012). A Markov decision process (MDP) is a POMDP where each state is fully observable and a Markov chain (MC) is an MDP with a single action (which can be omitted).
2.1 Policies and Expected Costs
A policy resolves the action choices in a POMDP, and is a function $\pi\colon\mathcal{H}\to\Delta(A)$ mapping histories to distributions over actions. Correspondingly, policies may also be belief-based $\pi\colon\mathcal{B}\to\Delta(A)$. The set of all history-based policies is $\Pi$. We seek to find a policy $\pi\in\Pi$ that minimizes the expected cost of reaching goal states $G\subseteq S$, also known as the stochastic shortest path (SSP) problem (Bertsekas,, 2005). For any trajectory $\omega$, the cumulative cost $\rho_{\lozenge G}\colon\Omega\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is (Forejt et al.,, 2011):
$\rho_{\lozenge G}(\omega)=\begin{cases}\infty&\forall t\in\mathbb{N},s_{t}\not\in G,\\ \sum_{t=0}^{\min\{t\;\mid\;s_{t}\in G\}-1}C(s_{t},a_{t})&\text{otherwise.}\end{cases}$ (1)
The SSP objective is to find an optimal policy $\pi\in\Pi$ that minimizes the expected cumulative cost $J^{\pi}_{T}$ of the trajectories generated by playing $\pi$ under the transition function $T$:
$\pi^{*}=\operatorname*{arginf}_{\pi\in\Pi}J^{\pi}_{T},\quad J^{\pi}_{T}=\mathbb{E}_{\pi,T}\bigg{[}\rho_{\lozenge G}\mid s_{0}\sim b_{0}\bigg{]}.$ (2)
In the fully observable setting of MDPs, dynamic programming can compute an optimal policy and the associated cost for the SSP problem. For POMDPs, finding optimal policies for the SSP problem is undecidable (Madani et al.,, 2003). Therefore, we approximate optimal policies with finite memory.
Finite-state controllers. A policy has finite-memory if it can be represented by a finite-state controller.
Definition 2 (FSC).
An FSC is a tuple $\pi_{f}=\braket{N,n_{0},\delta,\eta}$ where $N$ is a finite set of memory nodes and $n_{0}\in N$ the initial node, $\delta\colon N\times Z\to\Delta(A)$ is the stochastic action function, and $\eta\colon N\times Z\to N$ is the deterministic memory update. When $|N|=k$, the FSC is also called a $k$-FSC.
The performance of an FSC $\pi_{f}$ on a POMDP $M$ can be evaluated by computing the state-values $V^{\pi_{f}}\colon S\times N\to\mathbb{R}$ on the product Markov chain (Meuleau et al.,, 1999) via dynamic programming.
Recurrent neural networks. RNNs are (infinite) state machines parameterized by differential functions, which we use to learn information about the history in ^$\hat{\mathcal{H}}\subseteq\mathbb{R}^{d}$. The hidden size $d$ defines the length of the latent history vector. Analogously to an FSC, the RNN consists of a parameterized internal memory update $\hat{\eta}_{\phi}\colon\hat{\mathcal{H}}\times Z\to\hat{\mathcal{H}}$ that recurrently computes the new latent memory vector $\hat{h}\in\hat{\mathcal{H}}$ from the observations in a history $h\in\mathcal{H}$. Thus, the RNN represents a function $\text{RNN}\colon\mathcal{H}\to\hat{\mathcal{H}}$. When we append a fully connected layer with a $\operatorname*{softmax}$ activation function $\sigma^{\phi}\colon\hat{\mathcal{H}}\to\Delta(A)$ to the RNN, it yields an RNN policy network $\pi^{\phi}\colon\mathcal{H}\to\Delta(A)$ from histories to distributions over actions.
3 Robust Markov Models
A robust MDP (RMDP) is an MDP where the transition probabilities are not precisely given but belong to a predefined set of distributions known as the uncertainty set (Nilim and Ghaoui,, 2005). We fix a specific type of RMDP where the transition function maps to a probability interval or zero.
Definition 3 (RMDP).
An RMDP with interval uncertainty is a tuple $\mathcal{U}=\braket{S,A,\mathcal{T},C}$, where $\mathcal{T}\colon S\times A\to(S\to\mathbb{I}\cup\{0\})$ is the uncertain transition function that maps each transition to either a probability interval in $\mathbb{I}$, or probability 00 whenever the transition does not exist.
The uncertainty is resolved in a game-like manner where the agent (choosing the actions) and nature (choosing a probability distribution from the uncertainty set) alternate turns (Iyengar,, 2005). We assume independence between all state-action pairs, an assumption known as $(s,a)$-rectangularity, which means that nature’s choices are not restricted by any previous choice (Iyengar,, 2005; Wiesemann et al.,, 2013). Thus, the uncertainty set $\mathcal{T}$ is convex and factorizes over state-action pairs:
$\mathcal{T}=\bigotimes_{(s,a)\in S\times A}\mathcal{T}(s,a),\quad\mathcal{T}(s,a)=\bigg{\{}T(s,a)\in\Delta(S)\;\big{|}\;\forall_{s^{\prime}\in S}\colon T(s^{\prime}\mid s,a)\in\mathbb{I}\cup\{0\}\bigg{\}}.$ (3)
Furthermore, we assume that nature’s resolution of uncertainty is static. This means that, if the model contains a cyclic structure, nature’s choice does not change when state-action pairs are revisited.
RMDPs have two optimal value functions and associated optimal policies: one where the agent and nature play adversarially, and one where they play cooperatively. The former is known as the robust setting, and the latter as optimistic. For ease of presentation, we focus on the robust setting for the remainder of the paper. Optimal robust $\underline{\mathcal{Q}}$ and $\underline{\mathcal{V}}$ values can be computed by extending dynamic programming to robust dynamic programming (Iyengar,, 2005; Nilim and Ghaoui,, 2005). The optimal robust $\underline{\mathcal{V}}$ and $\underline{\mathcal{Q}}$ values are the fixed points $\underline{\mathcal{V}}^{*}$ and $\underline{\mathcal{Q}}^{*}$ of the following recursive equations111 To compute optimistic policies and values $\overline{\mathcal{V}}$ and $\overline{\mathcal{Q}}$, we only need to replace the inner supremum by an infemum. All our following results extend to the optimistic case by substituting the $\sup$ for an $\inf$. :
$\underline{\mathcal{V}}_{n}(s)=\min_{a\in A}\underline{\mathcal{Q}}_{n}(s,a),\quad\underline{\mathcal{Q}}_{n{+}1}(s,a)=C(s,a)+\sup_{T(s,a){\in}\mathcal{T}(s,a)}\!\left\{\sum_{s^{\prime}\in S}T(s^{\prime}\,|\,s,a)\underline{\mathcal{V}}_{n}(s^{\prime})\right\}.$ (4)
Under $(s,a)$-rectangularity and the interval uncertainty model, the inner optimization problem can be efficiently solved via a bisection algorithm (Nilim and Ghaoui,, 2005, Section 7.2). The robust state-values $\underline{\mathcal{V}}$ and optimal policy, which is memoryless deterministic under $(s,a)$-rectangularity, can be derived from the robust action values $\underline{\mathcal{Q}}^{*}$ in the same way as for standard MDPs (Iyengar,, 2005)
Robust POMDPs (RPOMDPs; Osogami,, 2015) extend POMDPs in a similar way that RMDPs extend MDPs by accounting for uncertainty in the transition and observation functions. Without loss of generality, we consider only uncertainty in the transition function, as any RPOMDP with an uncertain observation function can be transformed into an equivalent RPOMDP with deterministic observations (Bovy et al.,, 2024, Appendix B). Again, we focus on uncertainty given by intervals.
Definition 4 (RPOMDP).
An RPOMDP with interval uncertainty is a tuple $\mathcal{M}=\braket{S,A,\mathcal{T},C,Z,O}$, where $S,A,Z,O$, and $C$ are the sets of states, actions, and observations, observation function and the cost function from standard POMDPs, respectively. The uncertain transition function $\mathcal{T}\colon S\times A\to(S\to\mathbb{I}\cup\{0\})$ is the same as in RMDPs.
We again assume $(s,a)$-rectangularity and static uncertainty resolution. An instance of an RPOMDP $\mathcal{M}$ is a POMDP $M$ such that every transition probability of $T$ lies within its respective interval in $\mathcal{T}$. We also write $T\in\mathcal{T}$ and $M\in\mathcal{M}$ in that case. In RPOMDPs, the trajectories generated by the policy $\pi\in\Pi$ depend on the transition function $T\in\mathcal{T}$. Hence, the robust SSP objective becomes:
$\pi^{*}=\operatorname*{arginf}_{\pi\in\Pi}\mathcal{J}^{\pi}_{\mathcal{T}},\quad\mathcal{J}^{\pi}_{\mathcal{T}}=\sup_{T\in\mathcal{T}}J^{\pi}_{T}.$ (5)
Goal. Given an RPOMDP $\mathcal{M}=\braket{S,A,\mathcal{T},C,Z,O}$ with a set of goal states $G\subseteq S$, compute a policy $\pi\in\Pi$ in the form of a finite-state controller $\pi_{f}$ to minimize the worst-case expected cost $\mathcal{J}^{\pi}_{\mathcal{T}}$.
4 Supervised Learning of Robust FSCs
We now present our implementation of the PIP framework, named rFSCNet, to compute robust policies for RPOMDPs. We initialize with an arbitrary POMDP instance $M\in\mathcal{M}$ and a supervision policy $\pi_{M}$ that approximates the optimal policy $\pi^{*}$ for the POMDP $M$, which is used to train the RNN. Then, we enter the iterative procedure outlined below, which implements the two major parts of PIP, described in more detail in this section and the next section, respectively. Part one corresponds with the left-hand side in Figure 1, and consists of computing FSCs for adversarial POMDPs:
  1. 1.
    Simulate the supervision policy $\pi_{M}$ on $M$ and collect the histories and the action distributions of $\pi_{M}$ into a data set $\mathcal{D}$ (Section 4.1).
  2. 2.
    Train the RNN policy $\pi^{\phi}$ on the data $\mathcal{D}$ and extract an FSC $\pi_{f}$ (Sections 4.2 and 4.3).
Part two corresponds to the right-hand side of Figure 1. Here, we implement the evaluation of the FSC, and, subsequently, selection of the next adversarial POMDP with respect to the current policy:
  1. 3.
    Evaluate the FSC’s performance exactly through robust dynamic programming (Section 5.1).
  2. 4.
    Update the adversarial POMDP $M\in\mathcal{M}$ and supervision policy $\pi_{M}$ and start a new iteration at step 1 (Section 5.2).
We determine whether to stop the iteration at step (3) based on the FSC’s evaluation when the worst-case expected cost of the FSC satisfies a target threshold, or, alternatively, we may set a maximum number of iterations. When the procedure does not terminate, we continue from step 1 and save the best policy found thus far. We save the best policy based on the FSC’s evaluation.
4.1 Data Generation
To train the RNN, we generate a data set $\mathcal{D}$ by simulating the supervision policy $\pi_{M}\colon\mathcal{B}\to\Delta(A)$ on the current (adversarial) POMDP $M\in\mathcal{M}$. We discuss how to obtain $M$ and $\pi_{M}$ in the next sections. We aggregate a batch of $I\in\mathbb{N}$ finite-horizon histories and associated action distributions of the supervision policy $\pi_{M}$ up to length $H\in\mathbb{N}$.
During simulation for $i=1,\ldots,I$, we compute the beliefs $b_{t}^{(i)}$ from history $h^{(i)}_{t}$ derived from the probabilities of $M$ at each time step $0<t<H$, with $b_{0}$ from $\mathcal{M}$. Then, $\mu_{t}^{(i)}{=}\pi_{M}(b_{t}^{(i)})$ represents the action distributions of the supervision policy during the simulation, and $a^{(i)}_{t}\sim\mu^{(i)}_{t}$ is the action played in simulation $i$ at time $t$. Then, $\mathcal{D}=\{\{h_{t}^{(i)},\mu_{t}^{(i)}\}^{H}_{t=1}\}_{i=1}^{I}$ consists of $I$ simulations of $H$ histories and their associated action distributions.
Supervision Policies
As already mentioned in Section 2, computing an optimal policy in POMDPs for our objective is undecidable. Furthermore, our iterative procedure relies on computing several policies for different POMDPs. Thus, computational efficiency is a concern, and fast approximate methods are preferred. Therefore, we use an approximation $\hat{Q}_{M}$ of the belief-based values $Q_{M}\colon\mathcal{B}\times A\rightarrow\mathbb{R}$ for $M\in\mathcal{M}$. The supervision policy is then derived by taking the minimizing action, i.e., $\pi_{M}(b)=\operatorname*{argmin}_{a\in A}\hat{Q}_{M}(b,a)$. For training, we interpret these deterministic policies as stochastic by taking Dirac distributions. We consider the algorithms QMDP (Littman et al.,, 1995) and the fast-informed bound (FIB; Hauskrecht,, 2000) to compute the approximations QMDP and QFIB respectively.
QMDP may compute sub-optimal policies as it assumes full state observability after a single step, neglecting information-gathering (Littman et al.,, 1995). The values QFIB are tighter than the one given by QMDP since it factors in the observation of the next state. Although QFIB is computationally more expensive than QMDP, QFIB updates are still of polynomial complexity, allowing us to compute it for each $M\in\mathcal{M}$ efficiently, See Appendix B for more details.
Alternatives. We opt for QMDP and QFIB because of their computational efficiency. Nonetheless, other POMDP solution methods may also be used, such as Monte Carlo planning (Silver and Veness,, 2010) or variants of heuristic-search value iteration (Smith and Simmons,, 2004; Horák et al.,, 2018).
4.2 RNN Policy Training
We train the RNN policy $\pi^{\phi}$ on the data set $\mathcal{D}$ collected from $M\in\mathcal{M}$. The training objective $\Phi$ of the RNN is to minimize the distance between the distributions of the RNN policy $\pi^{\phi}$ and the distributions $\mu$ of the supervision policy $\pi_{M}$:
$\Phi=\min_{\phi}\frac{1}{|\mathcal{D}|H}\sum^{|\mathcal{D}|}_{i=0}\sum^{H}_{t=0}D_{\text{KL}}\big{(}\pi^{\phi}(h_{t}^{(i)})\parallel\mu_{t}^{(i)}\big{)},$
where $D_{\text{KL}}$ is the Kullback-Leibler divergence, the histories $h$ are inputs to the RNN, and the action distributions $\mu$ are labels. We optimize the parameters $\phi$ for this objective by calculating the gradient via backpropagation through time.
We opt for a model-based approach, where we use approximate solvers to compute the supervision policy $\pi_{M}$ for $M\in\mathcal{M}$ and train the RNN to imitate $\pi_{M}$ in a supervised manner. Thus, we know we can stop training once the loss for $\Phi$ is sufficiently low. Alternatively, one could employ a model-free objective such as recurrent policy gradients (Wierstra et al.,, 2007). However, the latter neglects the available information from the model and, therefore, requires a large number of samples to converge (Moerland et al.,, 2020). Furthermore, it is unclear when training should be stopped.
As the POMDP $M\in\mathcal{M}$ changes over the iterations to be adversarial against the current policy, we need a finite-memory representation of the policy to find the new adversarial POMDP $M^{\prime}$.
4.3 Extracting an FSC from an RNN
We cluster the hidden memory states of the RNN (Zeng et al.,, 1993; Omlin and Giles,, 1996) to find a finite-memory policy in the form of an FSC. There are multiple ways to achieve such a clustering. Prior work (Carr et al.,, 2021) uses a quantized bottleneck network (QBN; Koul et al.,, 2019) to reduce the possible memory states to a finite set. They train the QBN post hoc by minimizing the mean-squared error on the hidden states generated by the RNN. Alternatively, it can be trained end-to-end by updating the parameters with the gradient calculated with $\Phi$, which we name QRNN for quantized RNN. Moreover, similar to post hoc training of the QBN, we can run a clustering algorithm such as k-means++ (Arthur and Vassilvitskii,, 2007) to minimize the in-cluster variance of the hidden states. For post hoc training, we employ the histories in $\mathcal{D}$ to generate the RNN hidden states. We consider all three methods in this paper. For more details on the QBN, we refer to Section D.1.
Instead of through simulations, as done in Koul et al., (2019), we utilize the model to construct the FSC. The clustering determines a finite set $N$ from the RNN’s hidden states in $\mathcal{H}$. We find the FSC’s memory update $\eta$ by executing a forward pass of the RNN’s memory update $\hat{\eta}_{\phi}$ for each reconstruction of ${n}\in{N}$, which produces the next memory nodes ${n}^{\prime}\in N$ and RNN hidden state $\hat{h}\in\mathcal{H}$ for each $z\in Z$ by exploiting the batch dimension. Then, the action mapping $\delta$ for ${n}$ and $z$ is given by the distribution of the RNN policy network $\sigma^{\phi}(\hat{h})$ for the next memory state. The initial node $n_{0}$ is determined by the RNN’s initial state and we prune any unreachable nodes from the FSC.
5 Robust Policy Evaluation and Adversarial Selection of POMDPs
This section represents the right-hand side of Figure 1. For the FSCs found by the approach explained in the previous section, we present our methods for robust policy evaluation of the FSC and, subsequently, of selecting the worst-case adversarial POMDP $M^{\prime}\in\mathcal{M}$ from the uncertainty set.
5.1 Robust Policy Evaluation
The performance of a finite-memory policy represented by $\pi_{f}=\braket{N,n_{0},\eta,\delta}$ in an $(s,a)$-rectangular RPOMDP, such as we have, is given by the following robust value function, for all $\braket{s,n}\in S\times N$:
$\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n})=\sum_{a\in A}\Big{(}\delta(a\,|\,n,O(s))C(s,a)$
$\qquad+\sup_{T(s,a)\in\mathcal{T}(s,a)}\Big{\{}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\mid s,a)\delta(a\mid n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s^{\prime},n^{\prime}})\Big{\}}\Big{)}.$ (6)
The inner optimization problem is convex under $(s,a)$-rectangular and convex uncertainty sets, but solving it at each dynamic programming iteration requires $|S||A|$ linear programs. For computational tractability, we instead opt to evaluate the FSC on the RPOMDP via a product construction of a robust Markov chain, similar to the one used for evaluating FSCs in POMDPs (Meuleau et al.,, 1999).
Definition 5 (Robust policy evaluation).
Given an RPOMDP $\mathcal{M}=\braket{S,A,\mathcal{T},C,Z,O}$ with initial belief $b_{0}$ and an FSC $\pi_{f}=\braket{N,n_{0},\eta,\delta}$, the robust state-values of $\mathcal{M}$ under $F$ are given by the state-values in the robust Markov chain $\braket{S\times N,b^{\pi_{f}}_{0},\mathcal{T}^{\pi_{f}},C^{\pi_{f}}}$ where the state-space is the product of RPOMDP states $S$ and FSC memory nodes $N$, the initial state distribution is $b^{\pi_{f}}_{0}(\braket{s,n})=b_{0}(s)[n=n_{0}]$, and the robust transition and cost functions are:
$\mathcal{T}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n})$ $=\sum_{a\in A}\mathcal{T}(s,a)(s^{\prime})\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))],$
$C^{\pi_{f}}(\braket{s,n})$ $=\sum_{a\in A}\delta(a\,|\,n,O(s))C(s,a),$
where addition and multiplication over intervals follow the standard rules for interval arithmetic (Hickey et al.,, 2001). The robust value function $\underline{\mathcal{V}}^{\pi_{f}}\colon S\times N\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$, follows directly from robust dynamic programming on this robust Markov chain with the bisection algorithm, see Equation 4. Then, the robust value in the RPOMDP is given by $\sum_{\braket{s,n}\in S\times N}b_{0}^{\pi_{f}}(\braket{s,n})\underline{\mathcal{V}}^{\pi_{f}}(\braket{s,n})$.
We again assume full $(s,a)$-rectangularity on this product state-space, effectively leading to $(\braket{s,n},a)$-rectangularity. That is, nature’s choices are independent of the agent’s current memory, yielding a value $\underline{\mathcal{V}}^{\pi_{f}}$ that is a conservative bound compared to the value $\underline{\mathcal{V}}^{\pi_{f}}_{\star}$ where nature’s choices must be consistent with the agent’s memory. Appendix A contains the formal statement and proof. Using Definition 5, we can compute the performance of any FSC extracted from the RNN. Furthermore, the FSC representation and the robust state-values $\underline{\mathcal{V}}^{\pi_{f}}$ enable us to find the new associated adversarial POMDP $M^{\prime}\in\mathcal{M}$ to be used in the next iteration of our procedure.
5.2 Finding Adversarial POMDP Instances
We now construct a heuristic to find a new POMDP instance $M^{\prime}\in\mathcal{M}$ that constitutes the local worst-case instance for the current policy under $(s,a)$-rectangularity of the RPOMDP. Let $\pi_{f}=\braket{N,n_{0},\delta,\eta}$ be the current FSC. Given the robust value function $\underline{\mathcal{V}}^{\pi_{f}}$ computed at the previous step, we aim to find a POMDP $M^{\prime}\in\mathcal{M}$ that is adversarial to the FSC $\pi_{f}$ and, therefore, induces its worst-case value. We compute the following transition probability under the additional constraint that probabilities are dependent on the memory nodes $n\in N$ (i.e., $(s,a)$-rectangularity):
$\hat{T}(s,a)=\operatorname*{argsup}_{\hat{T}(s,a)\in\mathcal{T}(s,a)}\sum_{n\in N}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}\mathcal{T}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\mid\braket{s,n})\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}).$
We construct a linear program (LP) that precisely encodes our requirements. Let $\hat{T}(\cdot\,|\,s,a)$ be the vector of length $|S|$ of optimization variables representing the transition function probabilities for each state-action pair $(s,a)\in S\times A$. The LP is then given by:
$\max_{\hat{T}(s,a)}\quad$ $\sum_{n\in N}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}[n^{\prime}=\eta(n,O(s))]\;\delta(a\,|\,n,O(s))\hat{T}(s^{\prime}\mid s,a)\;\underline{\mathcal{V}}^{{\pi_{f}}}(\braket{s^{\prime},n^{\prime}})$
s.t. $\sum_{s^{\prime}\in S}\hat{T}(s^{\prime}\,|\,s,a)=1,\;\text{and,}\;\forall s^{\prime}\in S\colon\;\hat{T}(s^{\prime}\,|\,s,a)\in\mathcal{T}(s,a)(s^{\prime}).$
Solving this LP yields assignments for the variables that, for all $(s,a)\in S\times A$, satisfy $\hat{T}(s^{\prime}\,|\,s,a)\in\mathcal{T}(s,a)(s^{\prime})$ for all $s^{\prime}\in S$, and $\hat{T}(s,a)\in\Delta(S)$. By construction, the assignments are valid probability distributions inside each respective interval and yield a heuristic for the worst possible value for the given FSC under $(s,a)$-rectangularity. The variables $\hat{T}$ determine the transition function of the adversarial POMDP $M^{\prime}=\braket{S,A,\hat{T},C,Z,O}$. With this new POMDP instance $M^{\prime}\in\mathcal{M}$, we repeat the loop from the first step in Section 4.1 of our learning scheme until the FSC satisfies a target value or we reach the maximum number of iterations.
6 Experimental Evaluation
We empirically assess different aspects of rFSCNet according to the following research questions:
  1. 1.
    Robustness. Does rFSCNet provide robustness against model uncertainty? How does it compare to a Baseline that trains on a fixed POMDP, ignoring model uncertainty?
  2. 2.
    Configuration sensitivity. How do different configurations of supervision policies and clustering methods affect the performance of rFSCNet and Baseline?
  3. 3.
    Comparison with the state-of-the-art. How does rFSCNet’s performance compare to SCP (Cubuktepe et al.,, 2021), a state-of-the-art method for RPOMDPs?
  4. 4.
    Memory size sensitivity. How does the FSC size affect rFSCNet and SCP’s performance?
We use four well-known POMDP environments from the literature, extended to RPOMDPs: an Aircraft collision avoidance environment (Kochenderfer et al.,, 2015; Cubuktepe et al.,, 2021), and three partially observable grid-worlds that contain adversaries (Junges et al.,, 2021) named Avoid, Evade, and Intercept. We provide detailed descriptions and dimensions in Appendix E and run times in Appendix G. For full details on the tools and hyperparameters used, we refer to Appendix C.
Figure 2 compares the performance of rFSCNet to Baseline across $20$ seeds. Figure 3 shows rFSCNet’s median and minimum performance when configured with a memory size of $k\leq 9$, k-means++ clustering and QMDP as supervision policy, compared to the SCP method with two different FSC sizes $k\in\{3,9\}$. As the performance of SCP is not dependent on randomness, we only depict a single value. Other configurations can be found in Section F.1. The heatmaps showcase the effect of various memory sizes $k$ on the performance of both rFSCNet and SCP in Aircraft and Evade. Figure 5 compares rFSCNet across multiple supervision policy and clustering configurations to Baseline on Intercept. Figure 5 shows the difference in training performance, in terms of the RNN and QBN losses, between rFSCNet and Baseline on Intercept. Appendix F provides the complete set of results. Next, we address each research question based on these results.
Robustness. As seen in Figure 2, rFSCNet incurs lower expected costs on average than Baseline in two out of four environments. On Aircraft and Intercept, the results show a clear improvement in the performance of rFSCNet over Baseline. There, the minimum (best) value of Baseline across the seeds is worse than the median value of rFSCNet. On Evade, the baseline performs better, demonstrating that in this environment, it suffices to train on a single, fixed POMDP. On Avoid, rFSCNet does not improve over the median performance of Baseline, but the baseline may find worse policies, as seen from the outliers. Furthermore, Figure 5 shows that training on adversarial POMDPs, instead of a single fixed POMDP, is a more difficult learning target. Even though this makes the training task more challenging, rFSCNet still achieves good performance. Overall, we conclude that rFSCNet improves over the baseline in terms of performance on RPOMDPs.
Configuration sensitivity. Figure 5 depicts the performance of rFSCNet and Baseline across various configurations on Intercept. QRNN, which represents training the QBN end-to-end, proves less stable than post hoc, and k-means++ produces the best results. We do not observe a major difference between the use of QMDP and QFIB. All in all, it demonstrates that rFSCNet consistently outperforms Baseline across configurations, even while the configuration affects performance.
Figure 2: Boxplots depicting the minimum (i.e., best) robust value of the extracted FSC policies for both rFSCNet and Baseline across $20$ seeds, configured with QFIB and a QBN trained post hoc.
Robust values Aircraft Avoid Evade Intercept
RMDP (LB) $\underline{\mathcal{Q}}^{*}$ $94.24$ $18.05$ $31.19$ $16.99$
SCP $k=3$ $116.03$ $20.07$ $37.97$ 31.57
$k=9$ $116.58$ $29.51$ $39.78$ $101.12$
rFSCNet med. 103.30 18.51 37.61 $56.20$
min. 102.03 18.16 36.65 $34.95$
Figure 3: The robust values (lower is better) as computed by rFSCNet when compared to SCP. Left: median (med.) and minimum (min.) of the best values found across $20$ seeds. The lower bound (LB) is the optimal robust value $\underline{\mathcal{Q}}^{*}$ of the underlying RMDP. Bold indicates the best value found across methods. Italics indicate the best result apart from the minimum found by our method. Right: heatmaps comparing the median performance of rFSCNet to SCP under various memory sizes.
Comparison with the state-of-the-art. As seen in Figure 3, SCP performs well on Evade and best on Intercept when $k=3$. rFSCNet outperforms SCP on Aircraft, Avoid, and Evade by both the median and minimum performance across the $20$ seeds. When the memory size in SCP is set to $k=9$, which is also the highest memory size for rFSCNet we considered in these results, we outperform SCP across all benchmarks, both in terms of median and minimum performance.
Memory size sensitivity. The table of SCP across $k\in\{3,9\}$ in Figure 3 indicates that SCP performs worse with more memory, especially on Avoid and Intercept. The heatmaps of Figure 3 show that SCP also performs much worse on Aircraft and Evade when more memory is specified, as indicated by the high values. In contrast, rFSCNet is not sensitive to memory over-specification, exhibiting consistent performance across the memory sizes. These results demonstrate the benefit of learning the memory structure instead of specifying it beforehand, as done in SCP. We study these effects in more detail, including confidence estimates of the median performance of rFSCNet, in Section F.2.
Discussion and limitations. The results show that rFSCNet finds better robust policies compared to both SCP and Baseline methods, positively answering research questions (1), (2) and (3). To rFSCNet’s merit, our results also show that a learning-based approach through the RNN has the advantage of flexibility in determining the memory structure and sizes of the FSCs, positively answering research question (4). However, the performance is limited by the quality of the supervision policies we compute during training. For instance, this limitation could explain why rFSCNet performs worst on Intercept, as this benchmark relies on information gathering, an aspect on which QMDP is known to perform poorly on. Nonetheless, rFSCNet’s modularity allows for any POMDP policy to be applied in the RNN’s training procedure, allowing for trade-offs between quality and computational efficiency, depending on the task at hand.
Figure 4: Boxplots of the value of the best FSC across seeds on multiple configurations tested for rFSCNet and Baseline on Intercept.
Figure 5: The post hoc QBN and RNN training losses from the run with QMDP averaged over iterations on Intercept.
7 Related Work
Computing optimal robust policies for RMDPs has been studied extensively (Nilim and Ghaoui,, 2005; Iyengar,, 2005; Wiesemann et al.,, 2013). RMDPs are also used in model-based (robust) reinforcement learning (RL) to explicitly account for model uncertainty and enable efficient exploration (Jaksch et al.,, 2010; Petrik and Subramanian,, 2014; Suilen et al.,, 2022; Moos et al.,, 2022).
For RPOMDPs, early works extend value iteration and point-based methods to account for the additional uncertainty (Itoh and Nakamura,, 2007; Ni and Liu,, 2013; Osogami,, 2015), or use sampling over the uncertainty sets (Burns and Brock,, 2007). Ni and Liu, (2008) introduce a policy iteration algorithm for optimistic policies Chamie and Mostafa, (2018) consider robustifying a given POMDP policy to combat sensor imprecision. Nakao et al., (2021) extend value iteration for distributionally robust POMDPs, where the agent receives side information after a decision period, which is a less conservative setting. Extensions to value iteration for RPOMDPs typically do not scale well to the large state spaces (up to $13000+$) we consider in this paper. More recent methods compute FSCs for RPOMDPs via quadratic (Suilen et al.,, 2020) or sequential (Cubuktepe et al.,, 2021) convex programming. In contrast to our work, the convex optimization methods compute FSCs of a predefined size and structure. Additionally, they cannot compute FSCs in the optimistic case.
In RL, policy gradient algorithms incorporate memory using RNNs (Wierstra et al.,, 2007). Thereafter, RNNs serve as the baseline neural architecture for reinforcement learning in POMDPs (Ni et al.,, 2022). Recently, Wang et al., (2023) learned a stochastic recurrent memory representation for (non-robust) POMDPs using information from the model during training. RNNs have also successfully been used in a planning setting to compute FSCs for POMDPs (Carr et al.,, 2019, 2020, 2021). Robustness against model uncertainty has not yet been considered, and, to the best of our knowledge, no method exists yet that exploits the learning capabilities of RNNs in a robust planning setting.
8 Conclusion
We presented a novel framework accommodating supervised learning to train RNNs to act as robust policies for robust POMDPs. As part of our method, we extract FSCs from the trained RNN to evaluate the robust performance. Our experiments show that our approach of exploiting model-based information often yields more robust policies than the baseline approach with less information. Additionally, it is competitive with and even outperforms state-of-the-art solvers. In the future, we will investigate the generalization capabilities of RNN policies in a model-based setting, such as the ability to transfer robust policies to different and more complex domains.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning and, more specifically, robust sequential decision-making under partial observability. There are many potential societal consequences of our work due to the generality of RPOMDPs.
Acknowledgments and Disclosure of Funding
We would like to thank the anonymous reviewers for their valuable feedback, Eline M. Bovy and Merlijn Krale for their helpful discussions on this work, and Niels Neerhoff for allowing us to use the code of his master’s thesis. This research has been partially funded by the NWO grant NWA.1160.18.238 (PrimaVera) and the ERC Starting Grant 101077178 (DEUCE).
Appendix
Appendix A Robust Policy Evaluation
In this appendix, we show that robust policy evaluation (Definition 6) indeed provides a conservative upper bound to the actual robust Bellman equation Equation 6:
$\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n})=\sum_{a\in A}\Big{(}\delta(a\,|\,n,O(s))C(s,a)$
$\qquad+\sup_{T(s,a)\in\mathcal{T}(s,a)}\Big{\{}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\mid s,a)\delta(a\mid n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s^{\prime},n^{\prime}})\Big{\}}\Big{)}.$
Definition 6 (Robust policy evaluation (extended)).
Given an RPOMDP $\mathcal{M}=\braket{S,A,\mathcal{T},C,Z,O}$ with initial state distribution (belief) $b_{0}$ and an FSC $\pi_{f}=\braket{N,n_{0},\eta,\delta}$, the robust state-values of $\mathcal{M}$ under $F$, $\underline{V}^{\pi_{f}}\colon S\times N\to\mathbb{R}$, are given by the state-values in the robust Markov chain $\braket{S\times N,\mathcal{T}^{\pi_{f}},C^{\pi_{f}}}$ where the states are the product of RPOMDP states $S$ and FSC memory nodes $N$, and the robust transition and cost functions are:
$\mathcal{T}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n},a)$ $=\mathcal{T}(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]$
$\mathcal{T}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n})$ $=\sum_{a\in A}\mathcal{T}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n},a)$
$C^{\pi_{f}}(\braket{s,n})$ $=\sum_{a\in A}\delta(a\,|\,n,O(s))C(s,a),$
where addition and multiplication over intervals follow the standard rules for interval arithmetic (Hickey et al.,, 2001). Then, the robust state-values are defined as:
$\underline{\mathcal{V}}^{\pi_{f}}(\braket{s,n})$ $=C^{\pi_{f}}(\braket{s,n})+\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n})\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}).$ (7)
For an initial belief $b_{0}$ and initial memory node $n_{0}$, both state-based value functions can be extended to the value in the initial belief:
$\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{b_{0},n_{0}})=\sum_{s\in S}b_{0}(s)\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n_{0}}),$
$\underline{\mathcal{V}}^{\pi_{f}}(\braket{b_{0},n_{0}})=\sum_{s\in S}b_{0}(s)\underline{\mathcal{V}}^{\pi_{f}}(\braket{s,n_{0}}).$
Theorem 1.
Given the FSC $\pi_{f}$, the robust state-values $\underline{\mathcal{V}}^{\pi_{f}}$ of the robust Markov chain $\braket{S\times N,b_{0}^{\pi_{f}},T^{\pi_{f}},C^{\pi_{f}}}$ provide a (conservative) upper bound on the value $\underline{\mathcal{V}}^{\pi_{f}}_{\star}$ of $F$ under $(s,a)$-rectangularity in the RPOMDP with initial belief $b_{0}$. That is, $\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n})\leq\underline{\mathcal{V}}^{\pi_{f}}(\braket{s,n})$, and consequently $\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{b_{0},n_{0}})\leq\underline{\mathcal{V}}^{\pi_{f}}(\braket{b_{0},n_{0}})$.
Proof.
We show that $\underline{\mathcal{V}}^{\pi_{f}}(\braket{s,n})\geq\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n})$ for all $(\braket{s,n})\in S\times N$. Recall Equation 7. Omitting the constant $C^{\pi_{f}}(\braket{s,n})$, we rewrite the supremum as follows:
$\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\,|\,\braket{s,n})\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}})$
$=\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}\sum_{a\in A}T^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\mid\braket{s,n},a)\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}})$
$=\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}\sum_{a\in A}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}})$
$=\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{a\in A}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}).$
Under $(s,a)$-rectangularity of the RPOMDP, we can continue rewriting this supremum as:
$\sup_{T^{\pi_{f}}(\braket{s,n})\in\mathcal{T}^{\pi_{f}}(\braket{s,n})}\sum_{a\in A}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}})$
$=\sum_{a}\sup_{T^{\pi_{f}}(\braket{s,n},a)\in\mathcal{T}^{\pi_{f}}(\braket{s,n},a)}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}})$
$\geq\left(\sum_{a}\sup_{T(s,a)\in\mathcal{T}(s,a)}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\right).$
Inserting the constant $C^{\pi_{f}}(\braket{s,n})$ again, we derive:
$\sum_{a}\Bigg{(}\delta(a\mid n,O(s))C(s,a)\;\;+$
$\sup_{T(s,a)\in\mathcal{T}(s,a)}\sum_{s^{\prime}\in S}\sum_{n^{\prime}\in N}T(s^{\prime}\,|\,s,a)\delta(a\,|\,n,O(s))[n^{\prime}=\eta(n,O(s))]\underline{\mathcal{V}}^{\pi_{f}}(\braket{s^{\prime},n^{\prime}}\Bigg{)}=\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{s,n}).$
Since the inequality holds for each state-memory node pair, we also have for some initial belief $b_{0}$ and initial memory node $n_{0}$ that $\underline{\mathcal{V}}^{\pi_{f}}_{\star}(\braket{b_{0},n_{0}})\leq\underline{\mathcal{V}}^{\pi_{f}}(\braket{b_{0},n_{0}})$. ∎
Intuitively, the robust Markov chain, and thus its value function $\underline{\mathcal{V}}^{\pi_{f}}$, operates under $(\braket{s,n},a)$-rectangularity, meaning nature may choose a probability distribution for each state $s\in S$, memory node $n\in N$, and action $a\in A$ independently. In the RPOMDP, and thus the associated value function $\underline{\mathcal{V}}^{\pi_{f}}_{\star}$, nature operates under $(s,a)$-rectangularity, meaning it chooses probability distributions independently of the state $s$ and action $a$ but is restricted to choose the same probability distribution for each memory node $n$. The latter is more restrictive to nature, hence nature has fewer options to adversarially play against the agent. As a result, the agent’s cost may be lower than when nature’s choices depend on the agent’s memory.
This difference in semantics may also be explicitly encoded in a partially observable stochastic game by making the agent’s memory either observable or unobservable to nature (Bovy et al.,, 2024).
Appendix B Supervision Policies
This section elaborates on the POMDP approximations used for computing the supervision policies.
QMDP.
The QMDP algorithm (Littman et al.,, 1995) is an effective method to transform an optimal MDP policy to a POMDP policy by weighting the (optimal) action values $Q^{*}$ of the MDP to the current belief $b\in\mathcal{B}$ in the POMDP $M\in\mathcal{M}$:
$Q_{\text{MDP}}(b,a)=\sum_{s\in S}b(s)Q^{*}(s,a)=\sum_{s\in S}b(s)\left(C(s,a)+\sum_{s^{\prime}\in S}T(s^{\prime}\mid s,a)V^{*}_{\text{MDP}}(s^{\prime})\right),$
where $V^{*}_{\text{MDP}}$ is the optimal value of the MDP underlying the POMDP $M$.
Fast-informed bound.
The fast-informed bound (FIB; Hauskrecht,, 2000) is an approximation of the optimal value of the POMDP. It is tighter than the one given by QMDP since it includes a sum over the observation of the next state. The $Q$ values of FIB are defined as:
$Q_{\text{FIB}}(b,a)=\sum_{s\in S}b(s)\alpha^{a}(s)=\sum_{s\in S}b(s)\left(C(s,a)+\sum_{z\in Z}\min_{a^{\prime}\in A}\sum_{s^{\prime}\in S}T(s^{\prime}\mid s,a)[z=O(s^{\prime})]\alpha^{a^{\prime}}(s^{\prime})\right),$
where $\alpha^{a}\colon S\to\mathbb{R}$ for each $a\in A$ is a linear function, or alpha-vector, updated via:
$\alpha_{i+1}^{a}(s)=C(s,a)+\sum_{z\in Z}\max_{a^{\prime}}\sum_{s\in S}T(s^{\prime}\mid s,a)[z=O(s^{\prime})]\alpha_{i}^{a^{\prime}}(s^{\prime}).$
Appendix C Tools and Hyperparameters
We use the tools Storm Hensel et al., (2022) for parsing the models and PRISM Kwiatkowska et al., (2011) to compute the RMDP values for the lower bounds in Figure 3 and for robust policy evaluation. We build and train the RNN and the QBN using TensorFlow Abadi et al., (2016). The RNN cell is a gated recurrent unit (GRU, Cho et al.,, 2014). We initialize with a concrete POMDP instance $M\in\mathcal{M}$, where the intervals of the uncertainty sets are resolved to a value in the middle of the interval $[i,j]$ given by $\nicefrac{{i+j}}{{2}}$. For all the experiments, the simulation batch size is set to $I=256$, the maximum simulation length is set to $H=200$, and we run for a maximum of $50$ iterations. Both the RNN and QBN use an Adam optimizer (Kingma and Ba,, 2015) with a learning rate of $1\cdot 10^{-3}$. The hidden size of the RNNs was set to $d=16$. The experiments ran inside a Docker container on a Debian 12 machine. Our infrastructure includes an AMD Ryzen Threadripper PRO 5965WX machine with 512 GB of RAM. We train the neural networks on the CPU. The different seeds for the RNN-based methods were executed in parallel, each running on a single core. Multi-threading in the Gurobi LP solver (Gurobi Optimization, LLC,, 2023) used by SCP was enabled. In our initial tests, we considered hidden sizes $d\in\{3,16,64\}$, batch sizes $I\in\{128,256,512\}$, learning rates in the range of $[1\cdot 10^{-2},1\cdot 10^{-4}]$, and different number of iterations before arriving at our final values. We used the same infrastructure and experimental setup across methods.
Appendix D Network Architectures
In this section, we provide more details on the neural network architectures. Our post hoc QBN approach largely follows Koul et al., (2019) and Carr et al., (2021), apart from differences mentioned in Section 4.3. We used a batch size of $32$ for both networks during stochastic gradient descent.
D.1 QBN
Similar to prior work (Carr et al.,, 2021), we employ a quantized bottleneck network (QBN; Koul et al.,, 2019). It consists of an encoder $E\colon\hat{\mathcal{H}}\to[-1,1]^{l}$ that maps the output of the RNN to a latent encoding with $\tanh$ activation, where $l$ is the latent encoding dimension. The latent encoding is then quantized by a function $q\colon[-1,1]^{l}\to\beta^{B_{h}}$, where $\beta$ is the finite set of possible discrete values, for instance, $\beta=\{-1,0,1\}$ for three-level quantization. The bottleneck dimension $B_{h}$ is the number of quantized neurons. Lastly, there is a decoder $D\colon\beta^{B_{h}}\to\hat{\mathcal{H}}$ to reconstruct the input given the quantized encoding. The QBN represents a function $\mathcal{Q}\colon\hat{\mathcal{H}}\to\hat{\mathcal{H}}$ where $\mathcal{Q}(\hat{h})=D(q(E(\hat{h})))$ for all $\hat{h}\in\hat{\mathcal{H}}$. We train the QBN to minimize the reconstruction loss, i.e., mean-squared error, on the RNN’s memory representations derived from the histories in $\mathcal{D}$. The finite set of memory nodes extracted is formed by the Cartesian product ${N}=\times_{B_{h}}\beta$, and ${n}=q(E(\hat{h}))\in{N}$ is the discrete memory representation. Therefore, the extracted FSC’s memory size $|{N}|=|\beta|^{B_{h}}$ is directly controlled by $B_{h}$ and the quantization level, i.e., size of the set $\beta$. Note that the quantization level can be changed to be 2-level, i.e., with $B=\{-1,1\}$ using the $\operatorname{sign}$ function as $q$, resulting in different controller sizes.
To ensure the encoder $E$ maps to $[-1,1]$ we use $\tanh$ activation. The gradient of this activation function is close to one around the zero input. Thus, for the $3$-level quantization, we use a version $\tanh_{flat}$ of the $\tanh$ function in the encoder that is flatter around the zero input to allow for easier learning of quantization level 00, given by (Koul et al.,, 2019):
$\tanh_{flat}(x)=1.5\tanh(x)+0.5\tanh(-3x).$
To allow the gradient to pass through the quantization layer, we employ a simple straight-trough estimator that treats the quantization as an identity function during back-propagation (Li and Liu,, 2016). The quantization activation function was provided by the Larq library (Geiger and Team,, 2020). The encoder and decoder use a symmetrical architecture with $\tanh$ activation. The networks were quite small. The input and output sizes of the encoder and decoder were set to the hidden size $d$ of the RNN, with intermediate layers of sizes $8\cdot B_{h}$ and $4\cdot B_{h}$.
D.2 RNN
We use a Gated Recurrent Unit (GRU, Cho et al., (2014)) as the RNN architecture. Although there is no clear consensus between the Long Short-Term Memory (LSTM, Hochreiter and Schmidhuber, (1997)) architecture and the GRU, the latter has fewer parameters than the LSTM but does have the ability to learn long-term dependencies due to the forget gate. The forget gate is known to combat vanishing gradients that occur through the variant of stochastic gradient descent employed for sequential models, known as backpropagation through time. The inputs to the RNN were put through a learnable embedding layer. We trained the RNN policy using the method in section 4.2 with a categorical cross-entropy loss implemented in TensorFlow. To prevent exploding gradients in the RNN, we use a norm-clipped gradient and orthogonal weight initialization (Saxe et al.,, 2014) in the recurrent layer of the GRU, as recommended by Ni et al., (2022).
Table 1: Dimensions of each benchmark environment.
Instances Aircraft Avoid Evade Intercept
$|S|$ $13552$ $10225$ $4261$ $4802$
$|Z|$ $37$ $6968$ $2202$ $2062$
$|A|$ $5$ $4$ $5$ $4$
Appendix E Benchmark Descriptions
In this section of the Appendix, we describe the benchmarks studied in the paper. All environments are adapted with uncertain transition functions. The grid-world environments model the probability of taking multiple steps instead of a single one for each possible moving action, to which we assign the interval $[0.1,0.4]$. In Aircraft, we have two uncertainties: the probabilities of the pilot’s responsiveness and of the adversary changing direction, both mapping to the same $[0.6,0.8]$ interval. The dimensions of the benchmarks are given in Table 1. We specify the dimensions of the grid-worlds to the same sizes as set in Carr et al., (2023).
E.1 Aircraft Collision Avoidance
We consider a discretized and uncertain version of the aircraft collision avoidance problem (Kochenderfer,, 2015) as introduced in Cubuktepe et al., (2021).
Aircraft.
We follow the discretization procedure exactly and base our model on Cubuktepe et al., (2021), but slightly adapt it for our expected cost formulation. The objective is to minimize the expected cost, which models avoiding a collision with an intruder aircraft while taking into account partial observability (sensor errors) and uncertainty with respect to future paths. Crashes incur an additional cost of $100$ over the usual cost incurred of $1$ for each action. Furthermore, the only sink states are the goal states $G\subseteq S$.
E.2 Grid-worlds
We consider the grid-worlds introduced by Junges et al., (2021) but reformulate them as an expected cost (SSP) objective. All actions incur a cost of $c=1$, with an additional penalty of $c=100$ when in a bad state. Once again, the only sink states of the RPOMDP are the goal states $G\subseteq S$.
Avoid.
The Avoid benchmark models a scenario where a moving agent must keep a distance from patrolling adversaries that move with uncertain speed. Additionally, its sensor yields partial information about the position of the patrolling adversaries. The agent may exploit knowledge over the predefined routes of the patrolling adversaries.
Evade.
Evade is a scenario where a robot needs to reach a destination and evade a faster adversary. The agent has a limited range of vision but may scan the whole grid instead of moving, incurring the same cost as moving. A certain safe area is only accessible by the agent.
Intercept.
Intercept is the opposite of Evade because the agent aims to meet another robot before it leaves the grid via one of two available exits. Once the target robot has exited, the agent incurs an additional penalty of $c=100$ for each step before reaching a goal state. On top of the view radius, the agent observes a corridor in the center of the grid.
Appendix F Extended Experimental Evaluation
F.1 Configuration Study
Due to its modularity, rFSCNet allows for different configurations that may have an effect on its performance. In Table 2, we collect median and minimum results across different configurations of rFSCNet. The combination of QMDP and k-means++ proves best, which is what is shown in the table of Figure 3 in the paper. QRNN, the method that uses a QBN trained end-to-end, did not perform successfully on all environments. This is due to instability during training, caused by updating the QBN’s parameters with the gradients calculated from training the RNN, see Section 4.2. However, by directly encoding the clustering of the QBN into the RNN architecture during training, we observe an improvement in the median and minimum performance on the two successful runs, Avoid and Intercept, over training the QBN post hoc. We also show results for the Baseline when trained with the two different supervision policies.
rFSCNet Baseline
Clustering QRNN k-means++ QBN QBN QBN QBN
Supervision QMDP QMDP QMDP QFIB QMDP QFIB
Aircraft med $\times$ $103.30$ 102.95 $103.41$ $105.91$ $105.83$
min $\times$ $102.03$ $101.91$ 101.88 $104.66$ $104.60$
Avoid med $19.90$ 18.51 $20.19$ $19.43$ $18.83$ $18.70$
min $18.62$ 18.16 $18.57$ $18.53$ $18.39$ $18.35$
Evade med $\times$ $37.61$ $37.96$ $38.20$ 36.64 $36.67$
min $\times$ $36.65$ $36.98$ $37.07$ 36.06 $36.11$
Intercept med $66.00$ 56.20 $79.71$ $79.51$ $110.47$ $109.36$
min $42.92$ 34.95 $45.37$ $48.16$ $82.30$ $81.66$
Table 2: Evaluation across multiple configurations for rFSCNet and the Baseline. The values represent median (med.) and minimum (min.) robust values from the best FSCs computed of each run across 20 seeds. QRNN represents training the QBN end-to-end, see Section 4.3. $\times$ indicates a run failed. Bold indicates the best (med/min) performance for each environment, i.e., across the rows.
The full results in the form of boxplots are depicted in Figure 6.
Figure 6: Comparison of the robust values between rFSCNet and Baseline across configurations. For Avoid, we plot both with and without the run with QRNN, as it produces large outliers.
F.2 Extended Analysis on Various Memory Sizes
In this subsection, we study the memory comparison between rFSCNet and SCP in more detail. We chose Aircraft and Evade, as SCP appeared most consistent on these benchmarks. In this study, rFSCNet ran with a post hoc QBN and QMDP as supervision policy. We run for the maximal memory settings that we can restrict rFSCNet to when using the QBN, namely the values in the set $k\in\{3,4,8,9,16\}$, which is the size of the sets that can be found through binary $k\leq 2^{B_{h}}$ or ternary quantization $k\leq 3^{B_{h}}$, see also Section D.1. Figure 7 extends the right-side plot in Figure 3 with statistical details. Namely, we plot the standard deviation around the median values in the heatmap and show the global min and max of each method. We observe very stable performance for our method across the various memory sizes. Both on Aircraft and Evade, SCP shows relatively stable performance across memory sizes up to $k=9$. However, also on these benchmarks the performance drops when the required memory is set to a high level. rFSCNet evidently does not suffer from the same phenomenon. Furthermore, rFSCNet outperforms Aircraft on all memory settings and performs similarly or better than SCP on Evade.
Figure 7: Comparison of the robust values between rFSCNet and SCP. For rFSCNet, the error bars depict the standard deviation, and the dotted line shows for each method the global minimum and global maximum.
F.3 Loss Comparison
Figure 8 shows the RNN and QBN losses of the Baseline and rFSCNet on Aircraft and Avoid. Both runs employ a QBN and use FIB as the supervision policy. The QBN is trained individually from the RNN, i.e., post hoc. The results show that as the RNN loss decreases, the QBN reconstruction loss increases. This tells us that it gets increasingly hard to compress the hidden states of the RNN as they get more refined. An intuition is that the RNN learns to use a larger part of $\hat{\mathcal{H}}$ to represent the hidden states as training progresses, therefore making it harder to cluster the hidden states. Alternatively, one could train the QBN end-to-end. However, as we elaborate in Section F.1, this approach suffers from instability during training and, therefore, did not successfully run on all environments.
Figure 8: Comparison of RNN and QBN losses between rFSCNet and the Baseline over the iterations. The line shows the mean over 20 seeds and we plot the standard deviation around the mean for the RNN loss. On Aircraft and Evade, there is only a slight difference between the losses of the Baseline and rFSCNet. On the right, on Avoid, a big difference is visible.
Appendix G Run Times
For completeness, we report the run times of each procedure for every environment in Table 3. We report the run times using a QBN trained post hoc and FIB, which is the most expensive configuration in terms of computations. The RNN-based run times are averaged over the 20 seeds. We would like to note that the times given for SCP are user time and do not account for the total CPU time incurred by multi-threading. For the RNN-based approach, we see that the baseline is slightly faster in every environment except for Avoid, as it does not execute Step 4 of our method from Section 5.2, and does not need to recompute the supervision policies as the POMDP is fixed. The run times naturally increase for larger FSCs because the Markov chain used for robust policy evaluation grows larger when the FSC has more memory nodes. Our method spends the majority of its time in its robust evaluation, executed by robust value iteration (robust dynamic programming) in PRISM. Additionally, extraction from the QBN can take longer, as $|N|$ forward passes of the RNN are required. Typically, the worse the policies found, the longer it takes to perform robust dynamic programming. By comparing the run times of the baseline to our robust method, we see that the heuristic for finding worst-case instances does increase execution time. Finally, we would like to point out that the run times for SCP could be summed for a fair comparison, as running SCP for only $k=9$ yields much worse results than for $k=3$.
Table 3: Average run times in seconds across $20$ runs for rFSCNet and the Baseline, and the user time of the SCP method on each environment. Thus, both run-times represent a form of user time.
Robust ($k\leq 9$) Baseline ($k\leq 9$) SCP ($k=3$) SCP ($k=9$)
Aircraft $10562.51\pm 156.03$ $2518.39\pm 158.76$ $1133.8$ $2169.3$
Avoid $9987.82\pm 1209.82$ $12778.36\pm 1164.16$ $2167.9$ $6217.9$
Evade $5157.85\pm 131.58$ $1281.85\pm 58.73$ $872.7$ $3674.1$
Intercept $2501.95\pm 16.00$ $1624.33\pm 12.49$ $1884.9$ $3243.5$